OpenStudy (anonymous):

Write the complex number in polar form with argument (Ɵ) between 0-360 degrees. -πi

OpenStudy (anonymous):

lets try that again. absolute value is \[\pi\] angle is 270

OpenStudy (anonymous):

so polar form is \[\pi(\cos(270)+i\sin(270))\]